# The Devil Is in the Details: Tackling Unimodal Spurious Correlations for Generalizable Multimodal Reward Models

### v1

@@ 3.4 Shortcuts Hinder MM-RMs’ Generalization
(Shortcut-Failure Degradation). Given a standard MM-RM $\mathcal{M}^{e}$ trained on dataset $\mathcal{S}_{train}^{e}$, we employ the text-only RM $\mathcal{M}_{t}^{e}$ as a proxy for text-only shortcuts, which is initialized with the same model weights and trained on the identical dataset $\mathcal{S}_{train}^{e}$. This text-only RM is then used to split an o.o.d. test set $\mathcal{S}_{test}^{e^{\prime}}$ into two parts: one part is successfully classified by the text-only RM, termed the shortcut-success subset, while the other part, where the text-only RM fails, is termed the shortcut-fail subset. Shortcut-Failure Degradation refers to the difference in accuracy achieved by the standard MM-RM between the shortcut-success subset and the shortcut-fail subset.
@@ 3.5.1 Shortcut Opportunities Within Datasets
Data biases, or high-frequency co-occurrences of irrelevant features with labels, create opportunities for spurious correlations (Agrawal et al., 2018; Cadene et al., 2019). In reward modeling, length bias (Shen et al., 2023; Chen et al., 2024a) is the most intuitive example: In preference datasets, longer and well-structured responses are often preferred, leading to a learnable bias that diverges from human expectations.
@@ 3.5.2 Feature Acquisition Under Limited Representation Capacity
To verify this, we conduct investigation experiments using visual patch numbers as proxies for feature acquisition difficulty, with results presented and discussed in Appendix C. While results confirm their correlation with MM-RMs’ generalization, we note that simply scaling the visual patch numbers cannot serve as a universal solution for all data scenarios, and there exists a notable performance bottleneck, highlighting the importance of a systematic algorithm to address unimodal spurious correlations.
@@ 4.1 Training with Shortcut-Proxy
$\mathcal{L}_{sa}=-\mathop{\mathbb{E}}\limits_{({\boldsymbol{x}_{i}},y_{i})\in\mathcal{S}_{train}}[\operatorname{SFC}({\boldsymbol{x}_{i}},y_{i})\cdot\mathcal{L}({\boldsymbol{x}_{i}},y_{i})]$ (7)
Essentially, our shortcut-aware loss function utilizes the SFC value to dynamically reweight samples in training distribution: A sample with a higher SFC value indicates that the text-only branch struggles to model preferences, suggesting that multimodal fusion is crucial for robust learning, thus receiving an increased weight; conversely, a sample with a lower SFC value indicates that the text-only branch can easily discriminate it, leading to a reduced weight. We consider this reweighting mechanism as an adaptive method to shift the training data distribution towards environments where multimodal understanding is crucial. This approach helps prevent the MM-RM from becoming overly dependent on text-only shortcuts, which have been shown to significantly hinder the model’s generalization.
@@ 4.2 Shortcut-Disentangled Inference
Figure 8: Best-of-N downstream evaluation with N ranging from 1 to 64, using VLFeedback as the training set. Shortcut-aware MM-RM demonstrates better scalability, while standard MM-RM shows some degree of reward overoptimization (Gao et al., 2023). For any N less than 64, we use the unbiased estimator (Nakano et al., 2021) to calculate its average score, as detailed in Appendix D.
@@ 5 Experiments
5.1 Cross-Distribution Results
@@
5.2 Downstream Performance
To further assess the generalization capability of the Shortcut-aware MM-RM algorithm, we conduct evaluations on downstream tasks using a Best-of-N (BoN) selection strategy. The process involves generating 64 candidate responses from InternVL2-8B for each image-query pair. Multiple MM-RMs are then employed to score these candidates, and the response receiving the highest reward score is selected for downstream benchmark evaluation.
@@ 5.2 Downstream Performance
Table 2: Best-of-64 performance evaluation across downstream tasks, with GPT-4o serving as the universal judge. The first line refers to the results of InternVL2-8B without Best-of-N selection.
@@ Appendix B Investigating the Limited Role of Training Data Size in Generalization
Appendix C Investigating the Impact of Feature Acquisition Difficulty on Generalization
@@
Appendix D Best-of-N Unbiased Estimator

### v3

@@ 3.4 Shortcuts Hinder MM-RMs’ Generalization
(Shortcut-Failure Degradation). Given a standard MM-RM $\mathcal{M}^{e}$ trained on dataset $\mathcal{S}_{train}^{e}$, we employ the text-only RM $\mathcal{M}_{t}^{e}$ as a proxy for text-only shortcuts, which is initialized with the same model weights and trained on the identical dataset $\mathcal{S}_{train}^{e}$. This text-only RM is then used to split an o.o.d. test set $\mathcal{S}_{test}^{e^{\prime}}$ into two parts: one part is successfully classified by the text-only RM, termed the shortcut-success subset, while the other part, where the text-only RM fails, is termed the shortcut-fail subset. Shortcut-Failure Degradation refers to the difference in accuracy achieved by the standard MM-RM between the shortcut-success subset and the shortcut-fail subset.  [^1] We further analyze the implications of SFD in Appendix C.
@@ 3.5.1 Shortcut Opportunities Within Datasets
Data biases (Bahng et al., 2020; Jung et al., 2023), or high-frequency co-occurrences of irrelevant features with labels, create opportunities for spurious correlations (Agrawal et al., 2018; Cadene et al., 2019). In reward modeling, length bias (Shen et al., 2023; Chen et al., 2024a) is the most intuitive example: In preference datasets, longer and well-structured responses are often preferred, leading to a learnable bias that diverges from human expectations.
@@ 3.5.2 Feature Acquisition Under Limited Representation Capacity
To verify this, we conduct investigation experiments using visual patch numbers as proxies for feature acquisition difficulty, with results presented and discussed in Appendix D. While results confirm their correlation with MM-RMs’ generalization, we note that simply scaling the visual patch numbers cannot serve as a universal solution for all data scenarios, and there exists a notable performance bottleneck, highlighting the importance of a systematic algorithm to address unimodal spurious correlations.
@@ 4.1 Training with Shortcut-Proxy
$\mathcal{L}_{sa}=\mathop{\mathbb{E}}\limits_{({\boldsymbol{x}_{i}},y_{i})\in\mathcal{S}_{train}}[\operatorname{SFC}({\boldsymbol{x}_{i}},y_{i})\cdot\mathcal{L}({\boldsymbol{x}_{i}},y_{i})]$ (7)
Essentially, our shortcut-aware loss function utilizes the SFC value to dynamically reweight samples in training distribution: A sample with a higher SFC value indicates that the text-only branch struggles to model preferences, suggesting that multimodal fusion is crucial for robust learning, thus receiving an increased weight; conversely, a sample with a lower SFC value indicates that the text-only branch can easily discriminate it, leading to a reduced weight. We consider this reweighting mechanism as an adaptive method to shift the training data distribution towards environments where multimodal understanding is crucial. This approach helps prevent the MM-RM from becoming overly dependent on text-only shortcuts, which have been shown to significantly hinder the model’s generalization. We further explore the theoretical underpinnings of our proposed method in Appendix E.
@@ 4.2 Shortcut-Disentangled Inference
While our method explicitly targets text-only shortcuts, the framework’s modality-agnostic design allows for broader applicability to other forms of spurious correlations. We discuss this further in Appendix F.
@@
Figure 8: Best-of-N downstream evaluation with N ranging from 1 to 64, using VLFeedback as the training set. Shortcut-aware MM-RM demonstrates better scalability, while standard MM-RM shows some degree of reward overoptimization (Gao et al., 2023). For any N less than 64, we use the unbiased estimator (Nakano et al., 2021) to calculate its average score, as detailed in Appendix H.
@@ 5 Experiments
5.1 Cross-Distribution Evaluation
@@
Model Scale Analysis.
To further validate the robustness of our method across different model capacities, we extend experiments to 2B and 4B variants of InternVL2. As shown in Appendix G, the Shortcut-aware algorithm consistently improves generalization on o.o.d. scenarios across all model sizes, with average accuracy gains of +6.4 (2B) and +6.8 (4B) compared to standard MM-RMs.
5.2 Downstream Test-time Performance
We conduct evaluations on downstream tasks using a Best-of-N (BoN) test-time selection strategy. The process involves generating 64 candidate responses from InternVL2-8B for each image-query pair. MM-RMs are then employed to score these candidates, and the response receiving the highest reward score is selected for downstream evaluation.
@@ 5.2 Downstream Test-time Performance
Table 2: Best-of-64 evaluation across downstream tasks, with GPT-4o serving as the universal judge. The first line refers to the results of InternVL2-8B without Best-of-N selection.
@@ 7 Conclusion
Acknowledgements
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), Beijing Municipal Science and Technology Project (Nos. Z231100010323002), the Natural Science Foundation of China (No. 62306303).
@@ Appendix B Investigating the Limited Role of Training Data Size in Generalization
Appendix C Further Discussion on Shortcut-Failure Degradation
The Shortcut-Failure Degradation (SFD) metric serves as a specialized diagnostic tool for analyzing unimodal spurious correlations in MM-RMs. By measuring the accuracy gap between shortcut-success and shortcut-fail subsets, SFD directly quantifies a model’s dependence on text-only shortcuts. This makes it particularly valuable for identifying unimodal bias, as it isolates this specific failure mode from other potential robustness issues that might be captured by more general metrics.
Conceptually, SFD aligns with established robustness evaluation paradigms like worst-group accuracy (Sagawa et al., 2019; Nam et al., 2022; Chaudhuri et al., 2023), where the shortcut-fail subset represents our “worst group” in multimodal generalization. Both approaches focus on model performance in challenging scenarios where biased correlations break down. Just as worst-group accuracy reveals models that exploit majority-group features, SFD exposes models that over-rely on text shortcuts that may dominate training but fail during deployment.
From a practical standpoint, SFD’s significance becomes particularly clear in real-world applications. For instance, an MM-RM might learn to favor “long, descriptive captions” as a text shortcut during training, but would fail dramatically when deployed to domains where response quality critically depends on visual understanding (e.g., medical imaging or technical diagram interpretation). SFD precisely measures this vulnerability by evaluating performance when text shortcuts become unreliable, making it an essential metric for assessing MM-RM robustness in practical deployment scenarios.
Appendix D Investigating the Impact of Feature Acquisition Difficulty on Generalization
@@
Appendix E Further Discussion on the Proposed Method
In this section, we further discuss how our Shortcut-aware MM-RM learning algorithm provides both an information-theoretic justification for dynamic sample weighting and a practical connection to robust optimization, ensuring that MM-RMs prioritize learning truly multimodal reward functions.
Information-Theoretic Interpretation of SFC.
The Shortcut-Failure Coefficient (SFC) can be understood information-theoretically as quantifying the insufficiency of text-only shortcuts for prediction. Assuming optimized losses approximate conditional entropy terms, we can relate SFC to the underlying information structure:
$\operatorname{SFC}=\frac{\mathcal{L}_{t}}{\mathcal{L}_{t}+\mathcal{L}_{m}}\approx\frac{H(Y|T)}{H(Y|T)+H(Y|T,V)}$ (8)
Defining $\rho=\frac{I(V,Y|T)}{H(Y|T)}$ as the normalized contribution of visual information beyond text, we derive that $\text{SFC}\approx\frac{1}{2-\rho}$. This shows that SFC monotonically increases with $\rho$, meaning samples requiring stronger visual grounding receive higher weights. Thus, the Shortcut-Failure Coefficient provides a principled foundation for reweighting based on the unique predictive contribution of multimodal information.
Connection to Distributionally Robust Optimization.
Our method also shares conceptual connection with Group Distributionally Robust Optimization (Sagawa et al., 2019), , with each sample representing a group without explicit labels. We hypothesize that worst-case group probability depends on the degree to which samples cannot be predicted using text information alone, aligning with our observations of Shortcut-Failure Degradation.
By setting sample weights as: $w(\boldsymbol{x}_{i},y_{i})=\frac{\operatorname{SFC}(\boldsymbol{x}_{i},y_{i})}{\mathbb{E}_{(\boldsymbol{x}_{i},y_{i})}[\operatorname{SFC}(\boldsymbol{x}_{i},y_{i})]}$ , we effectively shift the training distribution from $\mathbb{P}$ to $\mathbb{Q}$, where: $\frac{d\mathbb{Q}}{d\mathbb{P}}(\boldsymbol{x}_{i},y_{i})\propto w(\boldsymbol{x}_{i},y_{i})$ These weights positively correlate with the optimal reweighting for minimizing worst-case risk, as they upweight samples where unimodal shortcuts are inadequate. This also aligns with our empirical observations in Section 5.1, where SFC-based reweighting significantly reduces generalization gaps.
Appendix F Adaption to Other Spurious Correlations
While our method explicitly targets text-only shortcuts, the framework’s modality-agnostic design allows for broader applicability to other forms of spurious correlations. To adapt the algorithm, two key modifications are required: proxy adaptation and dynamic reweighting. First, the auxiliary branch can be replaced with a problem-specific proxy (e.g., an image-only branch) to isolate the target spurious feature. Second, analogous to the Shortcut-Failure Coefficient (SFC) in Equation 5, failure signals from the new proxy can be computed to reweight training samples where the proxy struggles, thereby prioritizing cases requiring robust learning.
This flexibility stems from the core principle of our approach: dynamically shifting focus toward samples where spurious correlations fail, regardless of their modality. For instance, if background biases dominate a dataset (e.g., visual QA), training a background-only proxy and reweighting based on its failure rate would similarly encourage reliance on complementary information. The reweighting mechanism remains unchanged, ensuring the framework’s scalability while preserving its effectiveness in mitigating diverse spurious correlations.
Appendix G Model Size Effects on Cross-Distribution Performance
To investigate the impact of model capacity on generalization, we extend our experiments to 2B, 4B of the base architecture InternVL, as shown in Figure 10. While increasing model scale improves overall performance—with 8B achieving higher absolute accuracy than smaller models—we observe that scaling alone does not fully resolve unimodal shortcut reliance. For example, in challenging out-of-distribution scenarios (e.g., POVID → RLHF-V), all model scales exhibit poor performance, demonstrating that merely increasing model parameter scale is insufficient to completely address generalization issues.
Critically, our shortcut-aware algorithm consistently improves generalization across all model sizes, while demonstrating considerable scalability (see Figure 11). In out-of-distribution scenarios, the 2B, 4B, and 8B models achieve average accuracy gains of +6.4, +6.8, and +10.4 respectively, compared to standard MM-RMs.
This demonstrates that our dynamic reweighting method addresses fundamental limitations in multimodal alignment—complementing (rather than being superseded by) increased capacity. While scaling plays a certain role, explicit shortcut mitigation remains essential for robust reward modeling.
(a) InternVL2-2B
(b) InternVL2-4B
(c) InternVL2-8B
Figure 10: Cross-distribution evaluation of standard MM-RM across different model scales.
(a) InternVL2-2B
(b) InternVL2-4B
(c) InternVL2-8B
Figure 11: Cross-distribution evaluation of our Shortcut-aware MM-RM across different model scales.
Appendix H Best-of-N Unbiased Estimator